A project that provides annual cash flows of $18,000 for 6 years costs $61,000 today.

A. If the required return is 11 percent, what is the NPV for this project?

B. Determine the IRR for this project.

Show work please.
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=18000/1.11+18000/1.11^2+18000/1.11^3+18000/1.11^4+18000/1.11^5+18000/1.11^6

=76149.68

NPV=Present value of inflows-Present value of outflows

=76149.68-61000

=$15149.68(Approx)

b.Let irr be x%
At irr,present value of inflows=present value of outflows.

61000=18000/1.0x+18000/1.0x^2+18000/1.0x^3+18000/1.0x^4+18000/1.0x^5+18000/1.0x^6

Hence x=irr=19.24%(Approx).